OpenStudy (anonymous):

which method would be most efficient to solve the system? y=1/2x 2x+3y=28 a. graphing b. substitution c. elimination d. distributive

OpenStudy (anonymous):

i think its b

OpenStudy (anonymous):

correct.

OpenStudy (texaschic101):

thats what I would do....sub

OpenStudy (anonymous):

and what would the solution be a.(2,3) b. (8,4) c. (7,3.5) d. (2,14)

OpenStudy (texaschic101):

y = 1/2x sub that in for y in the other equation 2x + 3y = 28 2x + 3(1/2x) = 28 2x + 3/2x = 28 4/2x + 3/2x = 28 7/2x = 28 x = 28/(7/2) x = 28 * 2/7 x = 56/7 x = 8 y = 1/2x y = 1/2(8) y = 8/2 y = 4 answer is (8,4)

OpenStudy (anonymous):

for the following system of equations, what is the x-vaule of the solution? -x+2y=6 6y=x+18 a. -1 b. 0 c. 1 d. 6

OpenStudy (texaschic101):

-x + 2y = 6 -x = -2y + 6 -- multiply by -1 to make x positive x = 2y - 6 now sub this in for x in the other equation 6y = x + 18 6y = 2y - 6 + 18 6y - 2y = -6 + 18 4y = 12 y = 3 -x + 2y = 6 -x + 2(3) = 6 -x + 6 = 6 -x = 6 - 6 x = 0 x = 0